Find the equation to the tangent to the circle x(squared)+y(squared)=5 which is perpendicular to the line 4x-2y+3=0.

Grade:11

1 Answers

Ramesh V
14 years ago

slope of given line is 2, eqn is y = 2x + 3/2

its perpendicular eqn is of form , y = -1/2x + c where is c in intercept to be found ,as

this eqn is tangent , its perpendicular distance from center is equal to radius of 5 units

5 = c / (5/4)1/2

so , c = 5/2*(5)1/2

so eqn of tangent is : 2y + x - 5*(5)1/2 =0

--

regards

Ramesh